A manufacturing firm has one assembly line that is exhibiting the following workstation yields: WS1 = (0.98), WS2 =(0 .80), WS3 = (0.81), WS4 = (0.95), WS5 = (0.97) If workstations 2 and 3 were to increase their yields by just 5% (0.05), what would the percent increase in output be?
Answer #1

Current yield = 0.98*0.80*0.81*0.95*0.97 = 0.5852

New yield = 0.98*(0.8+0.05)*(0.81+0.05)*0.95*0.97 = 0.6601

Percent increase in output = (0.6601-0.5852)/0.5852

= 12.8 %